# A new expansion of the Riemann zeta function

### v1

@@ 3.6 An orthonormal expansion
then we have for $k\geq 0$
$X(k)=\int_{0}^{+\infty}e^{-t}\frac{t^{k}}{k!}\ f_{0}(t)dt$
By definition of Laguerre polynomials, we obtain for the coefficients $x_{m}$ the integral expression
@@
By Mellin-Parseval formula (cf. [T]), we have
$\frac{1}{\sqrt{2}}\int_{0}^{+\infty}{\varphi_{m}(t)}f_{0}(t)dt=\frac{1}{2\pi\sqrt{2}}\int_{-\infty}^{+\infty}\overline{\mathcal{M}(\varphi_{m})(\frac{1}{2}+it)}\ \mathcal{M}(f_{0})(\frac{1}{2}+it)\ dt$
Since (cf. [T]) we have $\mathcal{M}f(s)=\Gamma(s)\zeta(s)$ for $0<\Re(s)<1$, then, using (26), we get
$x_{m}=\frac{1}{2\pi}\int_{-\infty}^{+\infty}\overline{\Gamma(\frac{1}{2}+it)\frac{1}{2}P_{m}(\frac{1}{2}+it)}\Gamma((\frac{1}{2}+it))\frac{1}{2}\zeta(\frac{1}{2}+it)\ dt,$
since $Q_{m}(t)=\frac{1}{2}P_{m}(\frac{1}{2}+it)$, this is the integral expression (39).
@@
is also an orthonormal expansion in the space $\big(L^{2}(-\infty,+\infty),\frac{dt}{\cosh(\pi t)}\big)$.
Then we have
@@
We give next a formal evaluation of this last sum using Ramanujan’s summation.
3.7 A generating function
Let $\mathcal{L}$ the Laplace transform
$\mathcal{L}f(z)=\int_{0}^{+\infty}e^{-xz}f(x)\ dx=(e^{-xz},f).$
By Paley-Wiener theorem (cf. [Ho], [P]) , the transformation $\frac{1}{\sqrt{2\pi}}\mathcal{L}$ maps isometrically the space $L^{2}(]0,+\infty[,dx)$ to the Hardy space $H^{2}(\mathbb{C}_{+})$ of analytic functions $g$ in the half-plane $\mathbb{C}_{+}=\{z|\Re(z>0\}$, such that : there exists $M_{g}>0$ with
$\int_{\mathbb{R}}|g(x+iy)|^{2}\ dy\leq M_{g}\text{ for all }x>0.$
Each such $g$ has non-tangential limits $g(iy)=\lim_{x\to 0}g(x+iy)$ at almost every point of the imaginary axis. The space $H^{2}(\mathbb{C}_{+})$ is an Hilbert space with the inner product
$(g,h)_{H^{2}}=\int_{\mathbb{R}}\overline{g(iy)}{h(iy)}dy,$
and the functions $\psi_{m}$ $(m\geq 0)$, form an orthonormal basis of $H^{2}(\mathbb{C}_{+})$. If $f\in L^{2}(]0,+\infty[,dx)$. From
$e^{-xz}=\sqrt{2}\ \sum_{k=0}^{+\infty}\varphi_{k}(x)\psi_{k}(z).$
we get
$\frac{1}{\sqrt{2\pi}}\mathcal{L}\varphi_{m}(z)=\frac{1}{\sqrt{2\pi}}(e^{-xz},\varphi_{m})=\frac{1}{\sqrt{\pi}}\psi_{m}(z)$
The system $(\frac{1}{\sqrt{\pi}}\psi_{m})_{m\geq 0}$ is an orthonormal basis of $H^{2}(\mathbb{C}_{+})$. From the expansion
$f_{0}=\sum_{m\geq 0}(\varphi_{m},f_{0})\varphi_{m}$
we get the expansion
$\frac{1}{\sqrt{2\pi}}\mathcal{L}f_{0}(z)=\sum_{m\geq 0}(\varphi_{m},f_{0})\frac{1}{\sqrt{\pi}}\psi_{m}(z).$
and, by (38) we deduce
$\mathcal{L}f_{0}(z)=2\sum_{m\geq 0}x_{m}\psi_{m}(z).$
Since it is well known (cf. [G-R]) that
$\int_{0}^{+\infty}e^{-xz}f_{0}(x)dx=\log(z)-\Psi(z)-\frac{1}{z}$
then we get
@@ 3.7 A generating function
Equivalently this gives the generating function
$\sum_{m=0}^{+\infty}x_{m}u^{m}=\frac{1}{1-u}\Big(\log(\frac{1+u}{1-u})-\Psi(\frac{1+u}{1-u})\Big)-\frac{1}{1+u}.$ (42)
Remark 8.
@@
Then, with the results of examples 4-6, this relation gives
@@
Since $y_{m}=\int_{0}^{1}\psi_{m}(x)dx$ then by (41) we have
@@
Then (40) gives us the relation
$\int_{-\infty}^{+\infty}\Big|\zeta(\frac{1}{2}+it)\Big|^{2}\frac{dt}{\cosh(\pi t)}=2\log(2\pi)-2\gamma-1,$ (43)

### v2

@@ 3.6 An orthonormal expansion
it is well-known that $\mathcal{M}f_{0}(s)=\Gamma(s)\zeta(s)$ for $0<\Re(s)<1$. We easily verify that for $k\geq 0$ we have
$X(k)=\int_{0}^{+\infty}e^{-t}\frac{t^{k}}{k!}\ f_{0}(t)\ dt.$
Then by definition of Laguerre polynomials, we obtain
@@
This means that we have, in $L^{2}(]0,+\infty[)$, the orthonormal expansion
$f_{0}=\sqrt{2}\,\sum_{m=0}^{+\infty}x_{m}\,\varphi_{m}.$ (42)
@@
By (42) and Mellin-Parseval formula (cf. [T]), we have
$x_{m}=\frac{1}{\sqrt{2}}\int_{0}^{+\infty}{\varphi_{m}(t)}f_{0}(t)\ dt=\frac{1}{2\pi\sqrt{2}}\int_{-\infty}^{+\infty}\overline{\mathcal{M}(\varphi_{m})(\frac{1}{2}+it)}\ \mathcal{M}(f_{0})(\frac{1}{2}+it)\ dt,$
thus we get
$x_{m}$ $=$ $\frac{1}{2\pi}\int_{-\infty}^{+\infty}\overline{\Gamma(\frac{1}{2}+it)\frac{1}{2}P_{m}(\frac{1}{2}+it)}\Gamma((\frac{1}{2}+it))\frac{1}{2}\zeta(\frac{1}{2}+it)\ dt$
$=$ $\frac{1}{2\pi}\int_{-\infty}^{+\infty}\overline{\frac{1}{2}P_{m}(\frac{1}{2}+it)}\frac{1}{2}\zeta(\frac{1}{2}+it)\frac{dt}{\cosh(\pi t)}.$
Remark 8.
@@
is also an orthonormal expansion in the space $L^{2}\big(\mathbb{R},\frac{dt}{\cosh(\pi t)}\big)$. Then we have
@@
We now give a formal evaluation of this last sum using Ramanujan’s summation. Let $\mathcal{L}$ the Laplace transform
$\mathcal{L}f(z)=\int_{0}^{+\infty}e^{-xz}f(x)\ dx.$
We have (cf. [G-R]) for $Re(z)>0$, the Laplace transforms
$\mathcal{L}\varphi_{m}(z)$ $=$ ${\sqrt{2}}\ \psi_{m}(z),$
$\mathcal{L}f_{0}(z)$ $=$ $\log(z)-\Psi(z)-\frac{1}{z}.$
By Laplace transform of expansion (42) we get
@@
Then, with the evaluations of these sums given in lemma 3 and in (13), (14), (15), we get
@@
Since by (32) $y_{m}=\int_{0}^{1}\psi_{m}(x)\ dx$ then using (45) we have
@@
With (42) and (44) this gives
$\int_{0}^{+\infty}\Big(\frac{1}{e^{t}-1}-\frac{1}{t}\Big)^{2}\ dt=\frac{1}{2}\int_{-\infty}^{+\infty}\Big|\zeta(\frac{1}{2}+it)\Big|^{2}\frac{dt}{\cosh(\pi t)}=\log(2\pi)-\gamma-\frac{1}{2},$ (46)